One example of a nuclear fission reaction involving slowly moving neutrons is


1/0n + 235/92 U ---> 236/92 U* ---> 95/42 Mo + 139/57 La + (2) 1/0n


What is the total kinetic energy (in eV) of the products of the reaction? The relevant masses are:

235/92 U = 235.043 924 u; 95/42 Mo - 94.9058 u; 139/59 La = 138.9061 u, and 1/0n = 1.0087u


1
Expert's answer
2021-04-13T07:01:09-0400

To be given in question

Urenium mass =235.043"u"

Molebidnum mass =94.9058"u"

Neutron mass =1.0087"u"

Lanthanum mass =13 8.9061 "u"


To be asked in question

Kinetic energy in eV(election volt) =?

23592U +1n "\longrightarrow" 236U*"\longrightarrow" 9542 Mo+13957La+2(1n)

"M_{1}=M({235_{U}})+M_{{1}_{n}}"

"M_{2}=M_{{235}_{U}}+M_{{139}_{La}}+2M_{{1}_{n}}"

"\Delta M =M_{1}-M_{2}"

"M_{1}= 235.04+1.0087"

"M_{1}=236.0487 u"

"M_{2}=(94.9058+138.9061+2\times1.0087) u" "M_{2}=235.8293 u"

"\Delta M =M_{1}-M_{2}"

Put value

"\Delta M" =236.0487-235.8293 "u"

"\Delta M=0.2194 u"

"\Delta E=\Delta Mc^2"

"\Delta E=0.2194\times931.5"

"\Delta E=204.3711 mev"

1kg 235U fission energy release

"K.E=\frac{6.02\times 10^{23}}{0.235}\times{204.3711}"

"K.E=5.2353\times 10^{26}Mev""K.E=5.2353\times 10^{23}Kev"